OpenStudy (anonymous):

What degree Fahrenheit is it if the temperature is 15 degrees Celsius? Show all steps or reasoning. Be sure to check your answer for reasonableness.

OpenStudy (anonymous):

The formula to find change celsius to fahrenheit is: F=(C*(9/5))+32 F=(15*(9/5))+32 F=27+32=59 Hope it helps :)